OpenStudy (anonymous):

Assuming x > 0, what is the simplified fraction form for 15x times the square root of 3, all divided by the square root of 75x squared.

OpenStudy (anonymous):

Equation here:

OpenStudy (kira_yamato):

75 = 3 x 5^2 sqrt(75) = 5sqrt(3) sqrt(75x^2) = 5xsqrt(3) 15xsqrt(3) / 5xsqrt(3) = 3